# Not Just Change the Labels, Learn the Features: Watermarking Deep Neural Networks with Multi-View Data

### v1

@@ Abstract
With the increasing prevalence of Machine Learning as a Service (MLaaS) platforms, there is a growing focus on deep neural network (DNN) watermarking techniques. These methods are used to facilitate the verification of ownership for a target DNN model to protect intellectual property. One of the most widely employed watermarking techniques involves embedding a trigger set into the source model. Unfortunately, existing methodologies based on trigger sets are still susceptible to functionality-stealing attacks, potentially enabling adversaries to steal the functionality of the source model without a reliable means of verifying ownership. In this paper, we first introduce a novel perspective on trigger set-based watermarking methods from a feature learning perspective. Specifically, we demonstrate that by selecting data exhibiting multiple features, also referred to as multi-view data, it becomes feasible to effectively defend functionality stealing attacks. Based on this perspective, we introduce a novel watermarking technique based on Multi-view dATa, called MAT, for efficiently embedding watermarks within DNNs. This approach involves constructing a trigger set with multi-view data and incorporating a simple feature-based regularization method for training the source model. We validate our method across various benchmarks and demonstrate its efficacy in defending against model extraction attacks, surpassing relevant baselines by a significant margin.
@@ 1 Introduction
MLaaS empowers users to leverage the strong capabilities of DNNs, but it also exposes MLaaS providers to potential risks. Specifically, as these providers invest significant resources in developing the source model, protecting their intellectual property rights becomes a critical concern. Although the attackers do not have access to the source model nor the source training data, the attackers can still extract the DNN models’ functionality using black-box functionality stealing attacks [35, 30, 36, 42]. For example, model extraction attack leverages a surrogate model for imitating the output of the source model on a surrogate dataset for functionality stealing [35].
In order to enable the ownership verification of a stolen model, one of the most effective approaches is to use a trigger set for model watermarking [1, 46, 25, 33, 43, 18, 32, 2]. Specifically, during the training of the source model, the model owner utilizes a pre-selected trigger set, where the labels of the samples are known only to the owner. To verify ownership, the model owner assesses the output of the suspicious model. If the output aligns with the intended labels, the model owner can confidently assert ownership of the suspicious model.
@@
In this paper, we first introduce a novel perspective on trigger set-based watermarking methods based on feature learning. Specifically, we demonstrate the efficacy of employing multi-view data [3] as the trigger set. Multi-view data, defined as data exhibiting diverse features, is common in practice. For instance, a given horse image may possess the shape of a horse but mirroring the color of a dog, exhibiting two distinct features. In model extraction attacks, we demonstrate that when the source model utilizes color as a key feature for classifying a horse as a dog, the output of the source model on this multi-view trigger set can be successfully transferred to the surrogate model. This transfer makes the removal of the watermark a challenging task. Building on this perspective, we propose a simple approach to extract multi-view data from the source data for constructing the trigger set based on logit margin. To enhance the efficacy of using the trigger set, we additionally introduce a feature regularization loss, aiming to encourage the model to learn the desired features from this specific trigger set. Aligning with recent advancements in trigger-set-based watermarking methods [2, 20], we conduct experiments across various widely-used benchmarks. Our results demonstrate the superior performance of the proposed method compared to relevant baselines.
The contributions of the paper can be summarized as follows,
  • We present a novel perspective on trigger set-based watermarking methods, aiming to understand the conditions under which trigger set-based watermarking can effectively defend model extraction attacks. To the best of our knowledge, this perspective is novel in understanding trigger set-based watermarking methods.
  • We introduce an efficient trigger set-based watermarking approach, called MAT, by constructing a trigger set comprising multi-view data. The proposed method is simple to implement and does not require any changes to model architectures or the training process.
  • We evaluate MAT against various attack methods and the results demonstrate that MAT enables reliable ownership identification in situations where baseline methods fail.
@@ 2 Related Work
Watermarking Deep Neural Networks. Watermarking [14] was originally proposed to prove ownership of multimedia. The first model watermarking method [40] is proposed to protect intellectual property by adding watermarks to model parameters with a parameter regularizer during training. Based on this, More methods [4, 41] have been proposed to make model watermarks more difficult to detect and remove. However, these white-box model watermarking methods usually require access to the specific parameters of the suspect model to verify whether it contains a watermark or not. This is usually difficult to do in practical scenarios. Inspired by the backdoor attack technique [12], trigger-set based model watermarking methods [46, 1] are proposed to verify model ownership in black-box scenarios by simply querying the accuracy of a suspicious model on a specific trigger set. In order to make the watermark more robust, CEM [29] uses the trigger set constructed by conferrable adversarial examples when training the model, while margin-based method [20] uses adversarial training to maximize the margins of samples from the trigger set. The use of randomized smooth [5] during the training process has also been proven [2] to increase the non-removable watermark of the model.
Model Stealing. Although model watermarking is very effective in verifying ownership, it becomes vulnerable when adversaries employ attacks to attempt to remove the watermark. In the scenario of a white-box attack, the adversary can obtain the parameter weights of the model, so that some classic means of defending against backdoor attacks, such as fine-tuning and pruning [26], can easily remove the model watermark. In a more realistic case, black-box attacks such as distillation [17] and model extraction [35] can also steal clean models without watermarks through API queries of the source model. In particular, model extraction [35] is a functionality stealing method and is regarded as the most powerful black box attack currently [28, 20]. Some recent work [2, 20] claims to be resistant to functionality stealing, however under some restrictions. For example, random smoothing [2] requires limiting the size of perturbation for functionality stealing, while margin-based [20] requires the attacker to use the same training data set as the source model. Our approach effectively mitigates functionality theft in the most challenging and realistic scenarios. This includes situations where the attacker can employ an entirely different dataset, either mirroring the distribution of the source model training set or utilizing a completely unrelated dataset for model extraction.
@@ 3 Background
Threat model. Similar to existing trigger set-based watermarking methods [20, 2], we assume that adversaries aim to emulate the functionality of the source model using surrogate data, possibly from a distribution distinct from the source dataset used to train the model. They are aware of watermarking techniques but lack knowledge of the trigger set. Meanwhile, our defense operates in a black-box scenario, where defenders seek to establish ownership of the suspected model through a trigger set. Defenders have the capability to query predicted probabilities from the suspected model but may not have access to its parameters.
@@
where $D_{KL}$ is the Kullback-Leibler divergence [6] between two probability distributions. Note that the logits need to normalized via a softmax function. Although the attacker does not have access to the source dataset or source model parameters, the functionality of the source model can still be stolen based on the output of the source model on the surrogate data [35, 20].
Ownership Identification with a Trigger Set. To protect intellectual property, the owner of the source model must assess whether a particular surrogate model reproduces the functionality of the source model, a process commonly referred to as ownership identification [4, 37, 7, 43]. One of the most commonly used ownership identification method is based on watermarking using a trigger set [1, 18, 32, 2, 20]. In particular, for training the source model, the model owner randomly samples a subset of samples $\{(x_{k},y_{k})\}_{k=1}^{q}$ from the source dataset $S$ without replacement and replace the true label $y_{k}$ with $\hat{y}_{k}\neq y_{k}$ to construct a trigger set $D_{t}=\{(x_{k},\hat{y}_{k})\}_{k=1}^{q}$. Meanwhile, the trigger set is excluded from the original source dataset $S$ and the remaining clean training set is denoted as $D_{c}$ = $\{(x_{i},y_{i})\}_{i}^{m}$. With the additional trigger set, the model owner trains the source model by minimizing the following loss,
@@
Ideally, the model owner expects the surrogate model to produce the desired label for the trigger set. However, as outlined in Equation 2, since the attack is only based on the surrogate data for functionality stealing. As a result, there is no guarantee that the surrogate model will accurately extrapolate its predictions to the trigger set, especially considering the lack of correlation between the labels assigned to the trigger set and the images within it.
@@ 4 A New Perspective for Ownership Identification
In this paper, we introduce a novel perspective on the trigger set-based approach for watermarking based on feature learning. This viewpoint elucidates the conditions under which the surrogate model can potentially produce the desired label on the trigger set, leading to a new way for watermarking DNNs.
Conceptually, since the attacker does not rely on source data for extraction, transferring the prediction of the trigger set from the source model to the surrogate model can only based on the output of the source model on the surrogate data, a task that appears challenging. Here, we present a simple thought experiment to show the potential feasibility of such transfer. Suppose we have an image of a horse, where certain attributes, such as color, share a resemblance with those of a dog. Suppose we select this image as the trigger set and assign the label “dog” to this picture of a horse. During the training of the source model using Equation 8, the model is likely to rely on color cues in its classification process to identify an image as a dog. In this specific scenario, color serves as the exclusive distinguishing feature for classifying the image into the dog category. Now, in the context of a model extraction attack, let’s assume there exists a dog image within the surrogate dataset that shares the same color as the horse image in the trigger set. The source model, when presented with this dog image, would predict it as a dog, and this prediction would transfer to the surrogate model based on Equation 2. As a result, the surrogate model, having adopted this prediction, would subsequently categorize the horse image within the trigger set as a dog based on the color cue. To make the above idea concrete, we will leverage the following definition of multi-view data,
Multi-view data. The multi-view hypothesis [3] states that a given sample may exhibit multiple distinct features which can be used for classification. In the example above, the horse image exhibits both the distinctive shape characteristics of a horse and the color attributes of a dog. Suppose we have two classes: dog and horse, for simplicity. We assume each class $c$ has one feature $v_{c}\in\mathbb{R}^{p\times 1}$ which represent the mean feature of this class. We further assume the feature vectors of different classes are orthogonal, i.e, $v_{0}^{T}v_{1}=0$. For the sample $x^{c}_{i}$ belonging to class $c$, if the sample exhibit features from both classes, then we can represent the feature of this sample as $f_{i}^{c}=w^{0}_{i}v_{0}+w^{1}_{i}v_{1}$, where $w^{0}_{i}\in\mathbb{R}$ and $w^{1}_{i}\in\mathbb{R}$ are some constant weights. For example a given horse image may resemble a dog such that the feature of the horse can be written as $f_{horse}\approx 0.2v_{0}+0.8v_{1}$. As shown in [3], multi-view data commonly exists in real-world datasets, as natural images often possess diverse features that can be exploited for classification.
Watermarking using multi-view data. Here we show that multi-view data can be naturally used for constructing the trigger set. Assume that we assign a dog class to the horse image. Consider a simple binary classifier with weights $y=Wx+b$, where $W\in\mathbb{R}^{2\times p}$ and $b\in\mathbb{R}^{2}$. We use $W_{i}$ to denote the $i$-th row of $W$. Since the trigger set is trained together with clean data and the trigger set is relatively small compared with the clean data, then it is reasonable to assume the classifier can perfectly classify the two classes and the weights are perfectly aligned with the corresponding class, that is, $\cos(W_{i},v_{i})=1$. Given the sample $x^{1}_{i}$ with feature $f_{i}^{1}=w^{0}_{i}v_{0}+w^{1}_{i}v_{1}$, the logit can be computed as,
$\begin{split}z&=Wf_{i}^{1}+b=w^{0}_{i}*Wv_{0}+w^{1}_{i}*Wv_{1}+b\\ &=\begin{bmatrix}w^{0}_{i}W_{0}v_{0}+w^{1}_{i}W_{0}v_{1}+b\\ w^{0}_{i}W_{1}v_{0}+w^{1}_{i}W_{1}v_{1}+b\end{bmatrix}\end{split}$ (4)
$\text{softmax}(z)_{k}=\frac{e^{z_{k}}}{\sum_{j=1}^{2}e^{z_{j}}},\quad k=0,1$ (5)
If we assign a class 0 to this image, the binary cross-entropy loss can be simplified to $\ell=-\log(\text{softmax}(z)_{0})$. Based on our assumptions, $\text{softmax}(z)_{0}=\frac{e^{w^{0}_{i}W_{0}v_{0}+b}}{\sum_{j=1}^{2}e^{z_{j}}}$. Thus the source model will leverage the feature $v_{0}$ for classifying the images. In a model extraction attack, when a sample with feature $v_{0}$ is provided, the source model predicts class 0. This prediction is then leveraged to train a surrogate model, resulting in the surrogate model predicting the same label for the sample within the trigger set.
5 MAT
Based on above analysis, we introduce MAT, a novel watermarking method using trigger sets to defend model extraction attacks. MAT incorporates a novel trigger sample selection method and a new approach to source model training.
Bird (Airplane) Frog (Deer) Dog (Cat) Dog (Cat) Dog (Horse)
Figure 2: Some sample images in the trigger set selected by MAT on CIFAR10. The images exhibit features from different classes as expected. The true classes, along with the classes having the second-highest scores (in parentheses), are displayed.
Margin-based Trigger Set Selection. While our analysis in Section 4 demonstrates the effectiveness of transferring the source model’s predictions on the trigger set to the surrogate model through the utilization of data with features from different classes, identifying samples exhibiting multi-view features within a given source dataset remains unclear. To tackle this challenge, we introduce a simple and effective method for extracting multi-view data from the source dataset and subsequently employ them to construct the trigger set.
Given a source model $M_{\theta}$ and the logit $M_{\theta}(x)$ for a sample $x$ with label $y$. If the sample exhibit features from different classes, then the model will assign high scores to the respective classes. This intuition can be captured by the logit margin loss which is defined as $LM=\max_{j\neq y}M^{j}_{\theta}(x)-M^{y}_{\theta}(x)$. Given that the model makes the correction prediction, a large logit margin loss indicates that the features of sample $x$ comprise characteristics of both the true classes and the class with the second-highest score. Thus, we can select a trigger set consisting of samples with large logit margin loss. In practice, we first train the model on the source dataset $S$ using Equation 8, then select top $q$ samples with largest logit margin loss as the trigger set. The label of each selected sample is modified to the class $\hat{y}=\operatorname*{arg\,max}_{j\neq y}M^{j}_{\theta}(x)-M^{y}_{\theta}(x)$. Figure 2 shows the selected images using the proposed margin-based trigger set selection on CIFAR10. It can be observed that the images are difficult to classify and exhibit features from different classes.
Here we show a simple example with two classes: circle and square. With the logit margin loss, we would select the samples which are close to the decision boundary which exhibit both features of the circle class and the square class. Figure 1 a) shows the decision boundary of the source model on the source data. Figure 1 b) shows the decision boundary of the source model after training with the clean data and the chosen trigger set. It is noteworthy that this selection not only enhances watermarking performance but also impose minimal affect on clean accuracy, given that all other samples from the square class are considerably far from the chosen trigger set. This intuition is validated by the experimental results.
@@ 5 MAT
Feature Regularization. While the chosen trigger set possesses multi-view features, it is still important to encourage the source model to learn the desired features effectively. In the example shown in Figure 1 b), ideally, the source should employ the circle feature to classify samples in the trigger set. For instance, in a model extraction attack involving a circular image, the source model should be capable of transferring predictions on the trigger set to the surrogate model. Thus, we propose to enhance the model’s feature learning by incorporating a feature regularization loss. In particular, for a given sample $(x_{k},\hat{y}_{k})$ in the trigger set, we aims to push the feature of the sample to be close to the mean feature of class $\hat{y}_{k}$. This can be achieved by minimizing the following loss,
$\small\min_{\theta}\ell(\theta;D_{c})+\ell(\theta;D_{t})+\alpha\frac{1}{|D_{t}|}\sum_{(x_{k},\hat{y}_{k})\in D_{t}}\|f(x_{k})-f_{\hat{y}_{k}}\|_{2}.$ (6)
@@
where $\alpha$ is a balance parameter. As the mean feature of the class $\hat{y}_{k}$ is not directly available, we employ an approximation by computing them through the average features of samples belonging to the class $\hat{y}_{k}$. Specifically, we calculate the average features for each class during the previous epoch of model training and utilize these averages to regularize the feature learning of the trigger set in the current epoch. We will include the discussions and results of alternative regularization methods in the Supplementary.
Figure 1 c) demonstrates that additional feature regularization pushes trigger set features towards the circle class center, shifting the source model’s decision boundary accordingly. In Figure 1 d), the surrogate model’s decision boundary after a model extraction attack aligns with the source model’s predictions on the trigger set, as it aims to replicate them on the surrogate dataset.
@@
Crucially, MAT does not require the surrogate data to exhibit the features of the source dataset to successfully identify ownership. This is because the output of the source model for any given input can be used to approximate the decision boundary of the source model. By shaping the decision boundary of the source model using the chosen trigger set, the prediction of the source model on the trigger set can be transferred to the the surrogate model. We experimentally validate this in Sec. 6.4.
5.1 Ownership Verification with Hypothesis Test
@@
Our pipeline for ownership verification is similar to existing watermarking-based methods [24, 18, 20]. The defender can assert ownership of suspicious models by demonstrating statistically significant deviations in their behavior compared to a benign model. Specifically, we train a model $M^{c}_{\theta}$ directly on the clean dataset $D_{c}$, which serves as our benign model. We then evaluate the predictions of both the benign model $M^{c}_{\theta}$ and the suspicious model $\hat{M}_{\hat{\theta}}$ on the trigger set $D_{t}$, denoting their predictions as $P$ and $\hat{P}$, respectively. These predictions can be viewed as a binomial distribution, indicating whether watermarked data is classified as the target class [18]. Validation is conducted through a two-sample t-test, comparing the models’ predictions on the trigger set. If the resulting p-value is less than the significance level $\alpha$, defenders can claim ownership of the model.

### v2

@@ Abstract
With the increasing prevalence of Machine Learning as a Service (MLaaS) platforms, there is a growing focus on deep neural network (DNN) watermarking techniques. These methods are used to facilitate the verification of ownership for a target DNN model to protect intellectual property. One of the most widely employed watermarking techniques involves embedding a trigger set into the source model. Unfortunately, existing methodologies based on trigger sets are still susceptible to functionality-stealing attacks, potentially enabling adversaries to steal the functionality of the source model without a reliable means of verifying ownership. In this paper, we first introduce a novel perspective on trigger set-based watermarking methods from a feature learning perspective. Specifically, we demonstrate that by selecting data exhibiting multiple features, also referred to as multi-view data, it becomes feasible to effectively defend functionality stealing attacks. Based on this perspective, we introduce a novel watermarking technique based on Multi-view dATa, called MAT, for efficiently embedding watermarks within DNNs. This approach involves constructing a trigger set with multi-view data and incorporating a simple feature-based regularization method for training the source model. We validate our method across various benchmarks and demonstrate its efficacy in defending against model extraction attacks, surpassing relevant baselines by a significant margin. The code is available at: https://github.com/liyuxuan-github/MAT.
@@ 1 Introduction
MLaaS empowers users to leverage the strong capabilities of DNNs but also exposes MLaaS providers to potential risks. Specifically, as these providers invest significant resources in developing the source model, protecting their intellectual property rights becomes a critical concern. Although the attackers do not have access to the source model nor the source training data, the attackers can still extract the DNN models’ functionality using black-box functionality stealing attacks [39, 34, 40, 46]. For example, a model extraction attack leverages a surrogate model for imitating the output of the source model on a surrogate dataset for functionality stealing [39].
To enable the ownership verification of a stolen model, one of the most effective approaches is to use a trigger set for model watermarking [1, 50, 29, 37, 47, 19, 36, 2]. Specifically, during the training of the source model, the model owner utilizes a pre-selected trigger set, where the labels of the samples are known only to the owner. To verify ownership, the model owner assesses the output of the suspicious model. If the output aligns with the intended labels, the model owner can confidently assert ownership of the suspicious model.
@@
In this paper, we first introduce a novel perspective on trigger set-based watermarking methods based on feature learning. Specifically, we demonstrate the efficacy of employing multi-view data [3] as the trigger set. Multi-view data, defined as data exhibiting diverse features, is common in practice. For instance, a given horse image may possess the shape of a horse but mirror the color of a dog, exhibiting two distinct features. In model extraction attacks, we demonstrate that when the source model utilizes color as a key feature for classifying a horse as a dog, the output of the source model on this multi-view trigger set can be successfully transferred to the surrogate model. This transfer makes the removal of the watermark a challenging task. Building on this perspective, we propose a simple approach to extract multi-view data from the source data for constructing the trigger set based on logit margin. To enhance the efficacy of using the trigger set, we additionally introduce a feature regularization loss, aiming to encourage the model to learn the desired features from this specific trigger set. Aligning with recent advancements in trigger-set-based watermarking methods [2, 22], we conduct experiments across various widely used benchmarks. Our results demonstrate the superior performance of the proposed method compared to relevant baselines.
The contributions of the paper can be summarized as follows: 1) We present a novel perspective on trigger set-based watermarking methods, aiming to understand the conditions under which trigger set-based watermarking can effectively defend against model extraction attacks. To the best of our knowledge, this perspective is novel in understanding trigger set-based watermarking methods. 2) We introduce an efficient trigger set-based watermarking approach, called MAT, by constructing a trigger set comprising multi-view data. The proposed method is simple to implement and does not require any changes to model architectures or the training process. 3) We evaluate MAT against various attack methods and the results demonstrate that MAT enables reliable ownership identification in situations where baseline methods fail.
@@ 2 Related Work
Watermarking Deep Neural Networks. Watermarking [15] was originally proposed to prove ownership of multimedia. The first model watermarking method [44] is proposed to protect intellectual property by adding watermarks to model parameters with a parameter regularizer during training. Based on this, more methods [4, 45] have been proposed to make model watermarks more difficult to detect and remove. However, these white-box model watermarking methods usually require access to the specific parameters of the suspect model to verify whether it contains a watermark or not. This is usually difficult to do in practical scenarios. Inspired by the backdoor attack technique [12], trigger set-based model watermarking methods [50, 1] are proposed to verify model ownership in black-box scenarios by simply querying the accuracy of a suspicious model on a specific trigger set. To make the watermark more robust, CEM [33] uses the trigger set constructed by conferrable adversarial examples when training the model, while margin-based method [22] uses adversarial training to maximize the margins of samples from the trigger set. The use of randomized smooth [5] during the training process has also been proven [2] to increase the non-removable watermark of the model.
Model Stealing. Although model watermarking is very effective in verifying ownership, it becomes vulnerable when adversaries employ attacks to attempt to remove the watermark. In the scenario of a white-box attack, the adversary can obtain the parameter weights of the model, allowing classic defense mechanisms against backdoor attacks, such as fine-tuning and pruning [30], to easily eliminate the model watermark. In a more realistic case, black-box attacks such as distillation [18] and model extraction [39] can also steal clean models without watermarks through API queries of the source model. In particular, model extraction [39] is a functionality stealing method and is regarded as the most powerful black box attack currently [32, 22]. Some recent work [2, 22] claims to be resistant to functionality stealing, however under some restrictions. For example, random smoothing [2] requires limiting the size of perturbation for functionality stealing, while margin-based [22] requires the attacker to use the same training data set as the source model. Our approach effectively mitigates functionality theft in the most challenging and realistic scenarios. This includes situations where the attacker can employ an entirely different dataset, either mirroring the distribution of the source model training set or utilizing a completely unrelated dataset for model extraction.
@@ 3 Background
Threat model. Similar to existing trigger set-based watermarking methods [22, 2], we assume that adversaries aim to emulate the functionality of the source model using surrogate data, possibly from a distribution distinct from the source dataset used to train the model. They are aware of watermarking techniques but lack knowledge of the trigger set. Meanwhile, our defense operates in a black-box scenario, where defenders seek to establish ownership of the suspected model through a trigger set. Defenders can query predicted probabilities from the suspected model but may not have access to its parameters.
@@
where $D_{KL}$ is the Kullback-Leibler divergence [6] between two probability distributions. Note that the logits need to be normalized via a softmax function. Although the attacker does not have access to the source dataset or source model parameters, the functionality of the source model can still be stolen based on the output of the source model on the surrogate data [39, 22].
Ownership Identification with a Trigger Set. To protect intellectual property, the owner of the source model must assess whether a particular surrogate model reproduces the functionality of the source model, a process commonly referred to as ownership identification [4, 41, 7, 47]. One of the most commonly used ownership identification methods is based on watermarking using a trigger set [1, 19, 36, 2, 22]. In particular, for training the source model, the model owner randomly samples a subset of samples $\{(x_{k},y_{k})\}_{k=1}^{q}$ from the source dataset $S$ without replacement and replaces the true label $y_{k}$ with $\hat{y}_{k}\neq y_{k}$ to construct a trigger set $D_{t}=\{(x_{k},\hat{y}_{k})\}_{k=1}^{q}$. Meanwhile, the trigger set is excluded from the original source dataset $S$ and the remaining clean training set is denoted as $D_{c}$ = $\{(x_{i},y_{i})\}_{i}^{m}$, where q<m. With the additional trigger set, the model owner trains the source model by minimizing the following loss,
@@
Ideally, the model owner expects the surrogate model to produce the desired label for the trigger set. However, as outlined in Eq. 2, the attack is based solely on surrogate data for functionality stealing. Consequently, there is no guarantee that the surrogate model will accurately extrapolate its predictions to the trigger set, especially considering the potential lack of correlation between the assigned labels and the images within it.
@@ 4 A New Perspective for Ownership Identification
In this paper, we introduce a novel perspective on the trigger set-based approach for watermarking based on feature learning. This viewpoint elucidates the conditions under which the surrogate model can potentially produce the desired label on the trigger set, leading to a new way of watermarking DNNs.
Conceptually, because the attacker does not depend on the source data for extraction, transferring the predictions of the trigger set from the source model to the surrogate model can only rely on the output of the source model on the surrogate data—a task that seems challenging. Here, we present a simple thought experiment to demonstrate the potential feasibility of such a transfer. Suppose we have an image of a horse that shares certain attributes, such as color, with those of a dog. If we select this image as the trigger set and assign the label "dog" to it, we create a scenario where the attributes of a horse are mislabeled as those of a dog. During the training of the source model using Eq. 1, the model is likely to rely on color cues in its classification process to identify the image as a dog. In this specific scenario, color serves as the exclusive distinguishing feature for classifying the image into the dog category. Now, in the context of a model extraction attack, let’s assume there exists a dog image within the surrogate dataset that shares the same color as the horse image in the trigger set. The source model, when presented with this dog image, would predict it as a dog, and this prediction would transfer to the surrogate model based on Eq. 2. As a result, the surrogate model, having adopted this prediction, would subsequently categorize the horse image within the trigger set as a dog based on the color cue. To make the above idea concrete, we will leverage the following definition of multi-view data,
Multi-view data. The multi-view hypothesis [3] states that a given sample may exhibit multiple distinct features which can be used for classification. In the example above, the horse image exhibits both a horse’s distinctive shape characteristics and a dog’s color attributes. Suppose we have two classes: dog and horse, for simplicity. We assume each class $c$ has one feature $v_{c}\in\mathbb{R}^{p\times 1}$ which represents the mean feature of this class. We further assume the feature vectors of different classes are orthogonal, i.e., $v_{0}^{T}v_{1}=0$. For the sample $x^{c}_{i}$ belonging to class $c$, if the sample exhibits features from both classes, then we can represent the feature of this sample as $f_{i}^{c}=w^{0}_{i}v_{0}+w^{1}_{i}v_{1}$, where $w^{0}_{i},w^{1}_{i}\in\mathbb{R}$ are some constant weights. For example, a given horse image may resemble a dog such that the feature of the horse can be written as $f_{horse}\approx 0.2v_{0}+0.8v_{1}$. As shown in [3], multi-view data commonly exists in real-world datasets, as natural images often possess diverse features that can be exploited for classification.
Watermarking using multi-view data. Here we show that multi-view data can be naturally used for constructing the trigger set. Assume that we assign a dog class to the horse image. Consider a simple binary classifier with weights $y=Wx+b$, where $W\in\mathbb{R}^{2\times p}$ and $b\in\mathbb{R}^{2}$. We use $W_{i}$ to denote the $i$-th row of $W$. Since the trigger set is trained alongside the clean data and is relatively small in comparison, it is reasonable to assume that the classifier can perfectly classify the two classes and the weights are perfectly aligned with the corresponding class, that is, $\cos(W_{i},v_{i})=1$. Given the sample $x^{1}_{i}$ with feature $f_{i}^{1}=w^{0}_{i}v_{0}+w^{1}_{i}v_{1}$, the logit can be computed as,
$\begin{split}z&=Wf_{i}^{1}+b=w^{0}_{i}*Wv_{0}+w^{1}_{i}*Wv_{1}+b\\ &=\begin{bmatrix}w^{0}_{i}W_{0}v_{0}+w^{1}_{i}W_{0}v_{1}+b_{0}\\ w^{0}_{i}W_{1}v_{0}+w^{1}_{i}W_{1}v_{1}+b_{1}\end{bmatrix}\end{split}$ (4)
$\text{softmax}(z)_{k}=\frac{e^{z_{k}}}{\sum_{j=1}^{2}e^{z_{j}}},\quad k=0,1$ (5)
If we assign a class 0 to this image, the binary cross-entropy loss can be simplified to $\ell=-\log(\text{softmax}(z)_{0})$. Based on our assumptions, $\text{softmax}(z)_{0}=\frac{e^{w^{0}_{i}W_{0}v_{0}+b_{1}}}{\sum_{j=1}^{2}e^{z_{j}}}$ as $W_{0}v_{1}=0$. Thus the source model will leverage the feature $v_{0}$ for classifying the images. When a sample with feature $v_{0}$ is provided in a model extraction attack, the source model predicts class 0. This prediction is then leveraged to train a surrogate model, resulting in the surrogate model predicting the same label for the sample within the trigger set.
5 MAT
Based on the above analysis, we introduce MAT, a novel watermarking method using trigger sets to defend against model extraction attacks. MAT incorporates a novel trigger sample selection method and a new approach to source model training. In the Appendix, we illustrate our entire training pipeline.
Bird (Airplane) Frog (Deer) Dog (Cat) Dog (Cat) Dog (Horse)
Figure 1: Some sample images in the trigger set selected by MAT on CIFAR-10. The images exhibit features from different classes as expected. The true classes, along with the classes having the second-highest scores, are displayed in parentheses.
Margin-based Trigger Set Selection. While our analysis in Section 4 demonstrates the effectiveness of transferring the source model’s predictions on the trigger set to the surrogate model through the utilization of data with features from different classes, it remains unclear how to identify samples exhibiting multi-view features within a given source dataset. To tackle this challenge, we introduce a simple and effective method for extracting multi-view data from the source dataset and subsequently employ them to construct the trigger set.
Given a source model $M_{\theta}$ and the logits $M_{\theta}(x)$ for a sample $x$ with label $y$, if the sample exhibits features from different classes, then the model will assign high scores to the respective classes. This intuition can be captured by the logit margin loss which is defined as, $LM=\max_{j\neq y}M^{j}_{\theta}(x)-M^{y}_{\theta}(x)$. If the model makes the correct prediction, a large logit margin loss indicates that the features of sample $x$ comprise characteristics of both the true classes and the class with the second-highest score. Thus, we can select a trigger set consisting of samples with large logit margin loss. In practice, we first train the model on the source dataset $S$ using Eq. 1 and then select the top $q$ samples with the largest logit margin loss as the trigger set. The label of each selected sample is modified to the class $\hat{y}=\operatorname*{arg\,max}_{j\neq y}M^{j}_{\theta}(x)-M^{y}_{\theta}(x)$. Fig. 1 shows the selected images using the proposed margin-based trigger set selection on CIFAR-10. It can be observed that the images are difficult to classify and exhibit features from different classes.
In Fig. 2, we show a simple example with two classes: circle and rectangle. With the logit margin loss, we select samples close to the decision boundary that exhibit features of both the circle and rectangle classes. Fig. 2 a) shows the decision boundary of the source model on the source data. Fig. 2 b) shows the decision boundary of the source model after training with the clean data and the chosen trigger set. It is noteworthy that this selection not only enhances watermarking performance but also imposes minimal effect on clean accuracy, given that all other samples from the rectangle class are considerably far from the chosen trigger set. This intuition is validated by the experimental results.
@@ 5 MAT
Feature Regularization. While the chosen trigger set possesses multi-view features, it is still important to encourage the source model to learn the desired features effectively. In Fig. 2 b), ideally, the source should employ the circle feature to classify samples in the trigger set. For instance, in a model extraction attack involving a circular image, the source model should be capable of transferring predictions from the trigger set to the surrogate model. Thus, we propose to enhance the model’s feature learning by incorporating a feature regularization loss. In particular, for a given sample $(x_{k},\hat{y}_{k})$ in the trigger set, we aim to push the feature of the sample to be close to the mean feature of class $\hat{y}_{k}$. This can be achieved by minimizing the following loss,
$\small\min_{\theta}\ell(\theta;D_{c})+\ell(\theta;D_{t})+\alpha\frac{1}{|D_{t}|}\sum_{(x_{k},\hat{y}_{k})\in D_{t}}\|f(x_{k})-f_{\hat{y}_{k}}\|_{2}$ (6)
@@
where $\alpha$ is a balance parameter. As the mean feature of the class $\hat{y}_{k}$ is not directly available, we employ an approximation by computing them through the average features of samples belonging to the class $\hat{y}_{k}$. Specifically, we calculate the average features for each class during the previous epoch of model training and utilize these averages to regularize the feature learning of the trigger set in the current epoch. Fig. 2 c) demonstrates that additional feature regularization pushes trigger set features towards the circle class center, shifting the source model’s decision boundary accordingly. In Fig. 2 d), the surrogate model’s decision boundary after a model extraction attack aligns with the source model’s predictions on the trigger set, as it aims to replicate them on the surrogate dataset. We include the discussions and results of an alternative regularization method in the Appendix.
@@
Crucially, MAT does not require the surrogate data to exhibit the features of the source dataset to successfully identify ownership. This is because the output of the source model for any given input can be used to approximate the decision boundary of the source model. By shaping the decision boundary of the source model using the chosen trigger set, the prediction of the source model on the trigger set can be transferred to the surrogate model. We experimentally validate this in Sec. 6.4.
@@
Ownership Verification with Hypothesis Test. Our pipeline for ownership verification is similar to existing watermarking-based methods [26, 28, 13, 22]. The defender can assert ownership of suspicious models by demonstrating statistically significant deviations in their behavior compared to a benign model. Specifically, we train a model $M^{c}_{\theta}$ directly on the clean dataset $D_{c}$, which serves as our benign model. We then evaluate the predictions of both the benign model $M^{c}_{\theta}$ and the suspicious model $\hat{M}_{\hat{\theta}}$ on the trigger set $D_{t}$, denoting their predictions as $P$ and $\hat{P}$, respectively. These predictions can be represented as a binomial distribution, which indicates whether watermarked data is classified as the target class[19]. Validation is conducted through a two-sample t-test, comparing the models’ predictions on the trigger set. If the resulting p-value is less than the significance level, defenders can claim ownership of the model.